### v1

# MSC: A Marine Wildlife Video Dataset with Grounded Segmentation and Clip-Level Captioning

Abstract.
Marine videos present significant challenges for video understanding due to the dynamics of marine objects and the surrounding environment, camera motion, and the complexity of underwater scenes. Existing video captioning datasets, typically focused on generic or human-centric domains, often fail to generalize to the complexities of the marine environment and gain insights about marine life. To address these limitations, we propose a two-stage marine object-oriented video captioning pipeline. We introduce a comprehensive video understanding benchmark that leverages the triplets of video, text, and segmentation masks to facilitate visual grounding and captioning, leading to improved marine video understanding and analysis, and marine video generation. Additionally, we highlight the effectiveness of video splitting in order to detect salient object transitions in scene changes, which significantly enrich the semantics of captioning content. Our dataset and code have been released at https://msc.hkustvgd.com.
Marine datasets, marine videos, video captioning, referring video object segmentation.
Marine datasets, marine videos, video captioning, referring video object segmentation.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologiesccs: Computing methodologies Computer vision problemsccs: Computing methodologies Computer vision tasksccs: Computing methodologiesccs: Computing methodologies Computer vision problemsccs: Computing methodologies Computer vision tasks
Figure 1. The MSC dataset is recorded from 13 different countries, over 24.8 hours of marine video content. The dataset is associated with fine-grained annotations including clip-level textual descriptions provided by 18 biologists and pixel-level segmentation masks provided by 20 professionals.
Figure 2. Annotation Generation. The dataset curation consists of two tasks: creating segmentation masks and captioning. (a) We manually label marine videos to identify the target objects. We aim to build a fine-grained marine wildlife video-text dataset, performing the following stages: (b) Extracting clips from marine videos using annotated images. (c) Using a VLM to generate synthetic clip-level captions. (d) Refining the synthetic captions by domain experts. Please zoom-in for the best view.
1. Introduction
The recent advent of Artificial Intelligence (AI) has led to a new research capacity in life and environmental sciences, from on-ground animal biology (Stevens et al., 2024; Brookes et al., 2025; Tuia, 2025) to marine biology, e.g., coral detection (Zheng et al., 2024c), marine visual analysis (Zheng et al., 2024a; Mehrab et al., 2024).
The application of AI methods to ocean-related research is challenging due to the requirement of significant domain expertise and engineering work. Biologists have to collect data and manually label the data for specific marine species and tasks, then find and train a suitable model for each task. For real-world problems, off-the-shelf models often struggle to maintain comparative results. For example, visual grounding models, e.g., Grounding DINO (Liu et al., 2024b) and SAM2 (Ravi et al., 2024), are reliant on COCO’s predefined classes (Lin et al., 2014), and limited to user-defined input (e.g., customized text). Due to the presence of hundreds of marine species, these approaches are not applicable to the marine domain.
While Large Language Models (LLMs) can generate data at scale, they are prone to hallucination. Therefore, even with LLMs, human intervention is still needed to refine the synthetic data generated by the LLMs, especially in data-scarce domains. To address this issue, we propose a new marine video dataset, where the segmentation masks of marine objects are provided by annotators but the textual descriptions of the videos are generated automatically. Note that the MVK dataset introduced in (Truong et al., 2023) also used LLMs to generate image-level captions. Compared with MVK, our MSC provides more fine-grained and richer textual captions, generated from shorter temporal video segments (called clips) and validated by human experts. In addition, MSC is two times larger than MVK. Video Browser Showdown utilized MVK for known item search in (Vadicamo et al., 2024; Rossetto et al., 2024), finding MVK a challenging benchmark dataset for visual KIS tasks.
Natural questions arise: Why do we create a new video captioning dataset? and why do we focus on clip-level captioning? We argue that a clip-level video captioning dataset is essential for capturing richer semantic information, where each clip is defined as a semantically coherent segment. This leads to the development of our multimodal dataset comprising video-segmentation mask-text triplets, which explicitly link visual information with detailed textual descriptions and hold potential for multitask learning. Furthermore, we address the issue of hallucination prevalent in SOTA models, e.g., GPT-4.1 (gpt, T 41), suggesting that fine-grained, visually grounded data are crucial for generating accurate and reliable video captions.
In summary, our work makes the following contributions:
  • We propose MSC , the first real-world and large-scale marine video dataset, captured in various environments (see Figure 1).
  • We provide high-quality video-segmentation mask-text triplets for short-time video clips using a two-stage video annotation pipeline (see Figure 2).
  • We provide benchmark results on MSC across applications, including video captioning, plot/clip-level captioning, video generation, and visual grounding.
2. Related Work
Figure 3. Overview of the MSC dataset. (a) The dataset includes videos and clips, ranging from 1 to 30 clips per video. (b) and (c) Word cloud for noun and verb of marine captions generated by domain experts, respectively.
Sea floorReefsFishWrecksDiversPlants02,0002{,}0004,0004{,}0006,0006{,} [^0] Object count by category
[…]
00.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911HongKongEqyptWestPalmLajolleBigIslandTelAvivDublinTulambenSea floorReefsFishWrecksHuman diversAquatic plants
Figure 4. MSC Dataset. (a) The number of instances per category. (b) The average area of category. (c) Distribution of the number of instances by regions in the MSC dataset.
Underwater Video Datasets.
Marine videos have recently attracted considerable attention from the computer vision community. MarineInst (Zheng et al., 2024b) introduced a large-scale marine dataset with instance segmentation masks, facilitating image-level visual analysis tasks. Only 10% of the instance masks of MarineInst are annotated by humans while the remainder is generated by a segmentation model (e.g., SAM (Kirillov et al., 2023)). CoralSCOP (Zheng et al., 2024c) demonstrates the effectiveness of SAM in coral image segmentation, specifically addressing over-segmentation issues. SAM has been commonly used to create marine image instance segmentation datasets such as Watermask (Lian et al., 2023) and USIS10K (Lian et al., 2024).
Video-Text Datasets.
Recently, numerous large-scale video-text datasets, e.g., Koala-36M (Wang et al., 2024), HOIGen-1M (Liu et al., 2025), MiraData (Ju et al., 2024), have leveraged automatic captioning systems (e.g., GPT-4V) to generate video captions. MovieBench (Wu et al., 2024) provides video-, scene-, shot-level captions, also using GPT-4V, by incorporating additional semantic information. ViCaS (Athar et al., 2024) paves the way for new video-text datasets designed to evaluate video understanding tasks, including video captioning and visual grounding. Several datasets, e.g., BOVText (Wu et al., 2021), How2 (Sanabria et al., 2018), and VALUE (Li et al., 2021) are utilized for tasks such as video retrieval, captioning, video text spotting, and video question answering. Furthermore, domain-specific datasets, e.g., BASKET (Pan et al., 2025), are useful in applications such as classification and video generation. Although existing video-text datasets play an important role for training downstream tasks, e.g., video generation, they rely on LLMs that often produce hallucinated outcomes human interventions.
Language-guided Segmentation Datasets.
Early referring video object segmentation (RVOS) datasets, e.g., A2D Sentences (Gavrilyuk et al., 2018), J-HMDB Sentences (Gavrilyuk et al., 2018), DAVIS16-RVOS (Khoreva et al., 2019), DAVIS17-RVOS (Khoreva et al., 2019), and Refer-Youtube-VOS (Seo et al., 2020), focus on single-object segmentation, where language prompts are used to describe a single target object. These datasets often feature expressions, describing static attributes like color and shape. The number of expressions per object varies across the datasets. For instance, A2D Sentences and J-HMDB Sentences typically have 1-2 expressions per object, while DAVIS17-RVOS boasts an average of 7.5 expressions per object, and MeViS (Ding et al., 2023) provides 3.5 expressions per object. However, it is observed that there is a lack of video grounding datasets. To our knowledge, MSC is the first large-scale marine video text dataset for visual grounding. For visual grounding modeling, LLMs have been widely used to improve visual and language comprehension (Rasheed et al., 2024; He and Ding, 2024). Grounding DINO (Liu et al., 2024b), built upon Transformers, is designed to generate multiple bounding boxes as prompts for SAM.
Figure 5. Visualization of our web-based segmentation annotation tool.
Figure 6. The MSC ’s data organization.
3. Our MSC
3.1. Video Filtering
To choose marine videos for annotation, we considered three criteria (validated by human annotators): clarity, complexity, and diversity. (1) Clarity means that only videos whose objects can be seen with moderate clarity are selected. (2) Complexity means that videos containing objects in three or more distinct object types (e.g., fish, plants, reefs, divers) are selected. (3) Diversity means that we choose videos with distinct scenes or objects among those of the same diving site. Applying the above criteria, we finally selected 396 videos out of 2,743 videos.
3.2. Video Annotation
The annotation process involves two steps: creating video segmentation masks and captioning short-time clips. For the first step, annotators utilize a GUI tool to segment marine objects. We then use the manually annotated segmentation masks to identify target objects. In the second step, we provide high-quality clip-level captions using segmentation masks returned in the first step.
Step 1: Instance-level Video Segmentation.
We developed a web-based annotation tool for marine video object segmentation in Fig.  5. This annotation tool inherits from SAM (Kirillov et al., 2023) and (Wong et al., 2024) to produce high-quality pixel-wise segmentation masks (that we call pseudo-masks), and then allows annotators to refine the generated masks in an iterative manner. The annotators finally provide a category for each segmented mask. We focus on six categories, including fish, reefs, aquatic plants, wrecks, human divers, and sea floor.
Step 2: Captioning.
We leveraged LLMs to generating captions for our collected videos. However, we observed that the generated captions of long-time videos are often superficial because of the lack of detailed descriptions for the events included in the videos. To address this issue, we split long-time videos into short-time clips, each of which captures a single-shot event. We found that this step helps enhance the semantics of the generated captions.
After splitting long-time videos into short-time clips, we used GPT-4.1 (gpt, T 41), Gemini-2.0 Flash-Lite (gem, mini), Qwen-VL (Bai et al., 2023) to generate a textual description for each clip. The generated descriptions were then refined by biologists to reflect the semantic content specified by the segmented objects in the corresponding clips. The descriptions were refined to include the visual attributes and behavior (e.g., feeding, resting, breathing, social interactions, defense) of segmented objects and the background (surrounding environment) in the clips. Finally, the biologists produce a comprehensive caption for each video by aggregating the refined clip-level descriptions for the clips in that video, providing a concise summary of the video content.
Figure 7. Visualization of video-text segmentation mask triplets.
3.3. Analysis and Statistics
Diverse Lighting Conditions and Scenarios. Our MSC dataset was constructed by self-recording 71 visits across 20 distinct diving sites from 2011 to July 2024 (see Figure 1). Additionally, MSC is compiled with MVK (Truong et al., 2023), which encompasses recordings from various regions and under diverse illumination conditions. Data acquisition for this dataset was performed using GoPro cameras. Unlike common video datasets gathered through web crawling, our focus was specifically on marine organisms captured during dedicated diving expeditions. Importantly, a consistent data acquisition setup was maintained throughout the entire collection process. We show our data’s directory structure in Figure 6.
Table 1. Statistical overview of representative marine video datasets.
Dataset Total Duration (hours) Mean Duration (seconds) #Video #Country/Region
MVK (Truong et al., 2023) 12.3 29.9 1378 11
Our MSC 24.8 32.8 2743 20
Table 2. Video-level captioning on MSC dataset.
[…]
Qwen-VL-Chat 7B (Bai et al., 2023) 2023 0.0000 0.0759 0.1384 0.0148 0.0848
LLaVA 7B (Liu et al., 2024a) 2024 0.0125 0.1241 0.1798 0.0000 0.0689
PLLaVA 7B (Xu et al., 2024) 2024 0.0000 0.1114 0.1443 0.0028 0.0533
Gemini-2.0 Flash-Lite (gem, mini) 2025 0.0000 0.1251 0.1829 0.0571 0.0870
MovieBench (Wu et al., 2024) 2025 0.0000 0.1213 0.1790 0.0898 0.0914
Diverse Descriptions. We observed that textual descriptions obtained from LLMs pre-trained on large-scale datasets often suffer from hallucinations. To address this issue, we integrate the segmentation masks of target objects into textual descriptions, enabling the descriptions to be more focused on the target objects’ behaviors and the surrounding background, thereby eliminating hallucinations. This approach allows to create high-quality captions not only for visual grounding tasks but for the Text-to-Video (T2V) task via LLMs video captioning.
To further refine our synthetic captions, we engaged a team of 18 biologists to accurately identify marine organisms in the video footage. Synthetic captions generated by LLMs, i.e., GPT4.1 (gpt, T 41), QWen(Bai et al., 2023), Gemini 2.0 Flash-Lite (gem, mini), are employed to reduce the workload for manual captioning. Each short-time clip in our dataset was annotated with 1 to 4 text descriptions, comprising 3 synthetic and 1 human-written caption. The distribution of annotated videos is illustrated in Figure  3 (a). The word cloud of text descriptions is illustrated in Figure  3 (b).
Challenges with Object Scale and Imbalance. As shown in Figure 4 (a), while the numbers of fish and coral reef over 2,000 each, fish are predominantly small objects (0.67% of image area), whereas reefs are typically large 13.03%. Additionally, human divers are less frequent but consistently small objects (1.12% in Figure 4 (b)). Conversely, wrecks are rare but large objects (9.89% in Figure 4 (a, b)). This highlights an imbalance in both object quantity and scale in MSC dataset.
4. Challenges
Table 3. Clip-level captioning on MSC dataset.
[…]
Qwen-VL-Chat 7B (Bai et al., 2023) 2023 0.0000 0.0832 0.1541 0.0046 0.0696
LLaVA 7B (Liu et al., 2024a) 2023 0.0000 0.1337 0.1678 0.0057 0.0638
Gemini-2.0 Flash-Lite (gem, mini) 2025 0.0931 0.1736 0.2890 0.3679 0.1647
GPT4.1 (gpt, T 41) 2025 0.7196 0.5186 0.7844 4.9314 0.6139
Table 4. Visual Grounding performance on MSC dataset.
[…]
GroundingDINO + SAM2 2024 Labels 0.6543 0.7135
GroundingDINO + SAM2 2024 Caption 0.4447 0.5244
GLaMM + SAM2 2024 Caption 0.6727 0.7255
VideoGLaMM 2025 Caption 0.6812 0.7532
4.1. Video-level Captioning
Video captioning aims at generating a descriptive text for an input video sequence.
Baselines: In this work, we evaluated several prevailing visual-language models (VLMs) for video-level captioning in the MSC dataset. These models include Qwen-VL-Chat (Bai et al., 2023), LLaVA (Liu et al., 2024a), PLLaVA (Xu et al., 2024), Gemini (gem, mini), MovieBench (Wu et al., 2024). The models were used to generate captions at the video level for our marine video collection.
Qwen-VL uses Qwen as the language backbone and OpenCLIP ViT-bigG as the visual encoder, connected via a single-layer cross-attention module as the position-aware vision-language adapter. LLaVA employs a large language model, such as Vicuna, LLaMA, combined with a frozen vision encoder, such as CLIP ViT-L/14) via a single linear layer as a trainable projector. PLLaVA, an extension of LLaVA for videos, uses a parameter-free pooling strategy for video captioning. MovieBench is a data pipeline designed to create shot-level descriptions of scenes, with a particular emphasis on movie characters.
We set up to evaluate these VLMs with the guide prompt “Describe the video by following guidelines: you should give a paragraph with maximum 75 words; focus on the most obvious feature of the main objects in the initial frames; infer the behavior of the object (feeding, resting, breathing, social interactions, defense); and describe the background in about 10 words. Focus on fish, reefs, aquatic plants, wrecks, human divers, and sea floor. Omit the words ‘underwater’ and ‘shows’.”
[…]
Analysis: Gemini-2.0 and MovieBench perform well across the benchmarking metrics. Specifically, Gemini achieves 0.1251 (METEOR) and 0.1829 (ROUGE), while MovieBench achieves 0.0898 (CIDEr) and 0.0914 (SPICE). MovieBench is based on GPT-4.1, we observe that commercial models often produce comparative results.
4.2. Clip-level Captioning
Clip-level captioning aims to split a video into clips and then provide a caption for each clip.
Baselines: We also leverage various open-source and commercial VLMs like in the above video level captioning on our MSC dataset. With the similar setup to video captioning, we used the same prompt, but we input the frames from each clip rather than an entire video. We set the number of frames per clip to 10. To evaluate the clip-level captioning, we adopted the same metrics as in the Sec. 4.1.
Analysis: Regarding clip-level captioning benchmark, GPT-4.1 demonstrates superior performance, achieving the highest scores across all benchmark metrics. Gemini-2.0 ranks as the second-best performer in this comparison. Specifically, GPT-4.1 achieved scores of 0.7196 (BLEU), 0.5186 (METEOR), 0.7844 (ROUGE-L), 4.9314 (CIDEr), and 0.6139 (SPICE), while Gemini 2.0 yielded respective scores of 0.0931 (BLEU), 0.1736 (METEOR), 0.2890 (ROUGE-L), 0.3679 (CIDEr), and 0.1647 (SPICE). Note, we observe commercial models consistently achieve competitive performance in video and clip level captioning on this challenging domain in Table  3.
4.3. Visual Grounding
The visual grounding task in the MSC dataset involves linking a target model’s response to a user-specific text query by identifying marine creatures, objects, and their behaviors. This task requires the model to understand both spatial and temporal aspects in visually complex marine scenes characterized by high object variability, occlusion, and challenging lighting.
Given a caption {CAPTION} describing a clip, we prompt a target model using the format “{CAPTION}. Please respond with segmentation masks” to extract segmentation maps aligned with the textual description. The expected output highlights relevant spatial regions across frames, demonstrating the model’s ability to understand both spatial details (what appears in each frame) and temporal dynamics (how things change over time). We evaluated the grounding quality using mIoU and Recall metrics across annotated frames.
Baselines: We first assessed the performance of an open-vocabulary model, GroundingDINO (Liu et al., 2024b) combined with SAM2 (Ravi et al., 2024). For further validation, we benchmarked recent LLM-based visual grounding models, including VideoGLaMM (Munasinghe et al., 2024) and GLaMM (Rasheed et al., 2024). As GLaMM (Rasheed et al., 2024) is originally designed for image-based grounding, we extended it to the video domain by incorporating temporal capabilities through SAM2 (Ravi et al., 2024).
Analysis: While GroundingDINO + SAM2 show promise, their reliance on COCO-style categories limits their generalization ability to the marine domain. Performance notably degrades with natural, unconstrained captions and only improves when explicit label names are extracted for prompting. In contrast, LLM-based models better support spatio-temporal reasoning, establishing a more robust foundation for grounding in complex marine environments.
4.4. Text-to-video Generation
Our goal is to benchmark T2V models in generation of marine videos using clip-level captions from a video-text dataset. Pre-trained models were used in this experiment.
Baselines: We compared the open-source model Latte(Ma et al., 2025) and commercial models, i.e., Hailuo (hai, iluo), Kling 1.5 (kli, g 15), in marine video generation task. For the experimental setup, we used 50 clip-level caption prompts as input for the T2V models. Analysis: We observe commercial models often perform well on video generation across T2V metrics in Table 5. Specifically, Hailuo gains the superior result with 0.3236 (CLIP-T) while Kling 1.5 gains 0.997 on temperal consistancy metric. Pre-trained T2V models exhibit suboptimal performance in FID and FVD metrics due to insufficient diversity of underwater visual content in existing training datasets.
Table 5. Evaluation results with open-source model and commercial models in video generation.
[…]
Latte(Ma et al., 2025) 2025 0.3189 0.993 76.91 3123.05
Hailuo (hai, iluo) - 0.3236 0.9934 83.68 2007.37
Kling 1.5 (kli, g 15) 2024 0.3148 0.997 71.90 2820.24
5. Conclusion
This paper introduces MSC , the first large-scale video dataset of marine wildlife. The dataset contains fine-grained annotations, including object segmentation masks, clip-level captions, and video summaries. We developed an effective two-stage data annotation pipeline minimizing hallucinations by LLMs. Our associated benchmark includes video captioning, visual grounding, and text-to-video generation. We believe that MSC will contribute to facilitate research in marine video understanding.
Acknowledgements.
The work was partially supported by the Marine Conservation Enhancement Fund (MCEF20107 and MCEF22112), an internal grant from HKUST (R9429), HKUST Marine Robotics and Blue Economy Technology Grant. The authors would like to express their sincere gratitude to the “Sustainable Smart Campus as a Living Lab” (SSC) program at HKUST for its invaluable support. This program and staff not only provided essential funding and coordination, but also integrates sustainability into campus operations, which served as the living demonstration of the principles underpinning this research.